what is anthor way this can be written in .in distributive property? 2(m+n-q)=?
Answer by Mathtut(3670)

remember the distribution property of multiplication states:
:
a(b+c)=ab+ac.....in essence you are taking a and multiplying by each term in the parenthesis, no matter how many terms there are and adding the all together
:
so 2(m+n-q)=2m+2n-2q